# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

### v5

@@ 4.1 Meta-data Statistics of Reviews in PAMI
Figure 3: Statistics of the RiPAMI database. The samples in the RiPAMI database are characterized by a diversity of publication dates, scholar impact, reference numbers, etc., offering a comprehensive reflection of the state of reviews in the PAMI field.
@@
Number of the References The number of references in a survey paper may influence its credibility and reliability. As shown in 3 (b), the distribution of literature review references in the RiPAMI database follows a log-normal pattern.
Number of the Authors A review paper typically aims to provide a comprehensive overview of a specific topic. Involving multiple authors with diverse expertise could enhance the depth and breadth of the review. Fig. 3 (c) indicates that the majority of reviews are written by fewer than 10 authors.
Number of the Citations We count the citations of papers in RiPAMI and plot them in Fig. 3 (d). A power law distribution of received citations could be found. This phenomenon where a small subset of papers receives the majority of citations is sometimes referred to as the “Matthew Effect” or the “Pareto principle”, as reported in [83, 18].
Length of Reviews The words count and pages number of a review article can partially reflect its depth and breadth. In general, more detailed and comprehensive reviews tend to contain a higher word count and longer pages, thus requiring more time and effort to conduct and comprehend. As shown in Fig. 3 (e) and (f), review articles in the PAMI field typically range from 5,000 to 20,000 words and span 10 to 40 pages. At an average reading speed of 240 words per minute [17], reading a full review would take over 20 minutes.
@@
Further statistical details, such as maximum (Max), minimum (Min), mean, median, and mode, are available in Tab. II.
TABLE II: Statistical Meta-data Summary of RiPAMI
@@
4.2 Review Feature Statistics
@@
As discussed in Sec. 3, review articles in the PAMI field exhibit shared features. In this subsection, we apply information extraction techniques to analyze the review articles within the RiPAMI database, aiming to explore the overarching statistical patterns and trends within PAMI reviews. Specifically, we introduce an ensemble query approach based on LLMs to extract key information from the raw content, determining whether a review (1) Taxonomy: explicitly outlines a taxonomy for existing methods; (2) PRISMA: includes a dedicated section specifying inclusion/exclusion criteria, similar to PRISMA guidelines [92]; (3) Preliminary: provides background knowledge in a distinct section; (4) Benchmark: presents tables or figures comparing quantitative performance across methods; (5) Application: discusses current applications within the field; (6) Discussion: includes a section addressing challenges, future directions, or limitations of the field. These six features are recorded in binary format (0/1) and are used for subsequent statistical analysis. Details regarding the extraction of these content features are provided in Appendix B.
Figure 4 presents our analysis of content features across a sample of over 3,000 review articles. We found that the majority of reviews tend to include a dedicated section discussing challenges, future directions, or opportunities within the investigated field. The high prevalence of discussion of future and existing challenges reflects established academic norms, where reviews are expected to identify unresolved issues and propose future research paths. Notably, the proportion of reviews adhering to PRISMA in the PAMI field remains below 20%, regardless of whether they have been peer-reviewed or not. This indicates that there remains considerable scope for the advancement of rigorous systematic reviews within the PAMI field.
@@
Figure 4: Statistical Analysis of Review Features in the RiPAMI Database Using LLM-based Information Extraction Techniques: “Published” refers to peer-reviewed review papers, while “Non-published” denotes preprints without peer review.
@@
Figure 5 illustrates the changing proportions of six key review content features over the years. Most of the analyzed features show a steady upward trend over time, with “Discussion” and “Preliminary” seeing the most notable increases. The proportion of “PRISMA”, which reflects the methodological rigor of reviews, has risen by over 10% between 2015 and 2024, indicating that reviews in the PAMI field are becoming more structured and methodologically sound.
4.3 The Importance of References
As the saying goes, ’You can’t make something out of nothing.’ The same principle applies to literature reviews: a well-crafted review is typically built upon a thorough analysis of high-quality references. To further explore the relationship between the quality of references in literature reviews and the review’s impact, we conduct a visual analysis of the distributions of $ARQ$ (representing the quality of references), $S_{MP}$ (representing the recency of references), and the corresponding $TNCSI$ (representing the scholar impact) for reviews published after 2020, as shown in Fig. 6.
Figure 6: Visualization of references quality of human-authored and AI-generated literature reviews. Each point represents a review sample. The $TNCSI$ exhibits a positive correlation with $ARQ$ and $RQM$, and a negative correlation with $S_{MP}$.
TABLE III: Correlation between Reference Quality and TNCSI. “Sig.” indicates significance, and ✓indicates the corresponding p-value less than 0.05.
Metric Pearson Pearson p-value Spearman Spearman p-value
$ARQ$ 0.43 $6.99\times 10^{-40}$ 0.42 $1.11\times 10^{-37}$
$S_{mp}$0.27 $6.29\times 10^{-16}$0.32 $2.21\times 10^{-21}$
$RQM$ 0.53 $1.34\times 10^{-62}$ 0.51 $1.86\times 10^{-57}$
To further quantify the relationship between reference features and cumulative academic impact, we present the Pearson and Spearman correlation coefficients between various indicators and $TNCSI$ in Tab. III. Both $ARQ$ and $RQM$ show moderate positive correlations with $TNCSI$, with Pearson and Spearman coefficients around 0.43 and 0.53, respectively. These correlations are highly significant, as indicated by p-values close to zero. In contrast, $S_{mp}$ exhibits a negative correlation, as indicated by Pearson and Spearman coefficients of -0.27 and -0.32, respectively. These statistically validated results demonstrate the importance of reference quality in influencing academic impact of reviews.
4.4 Insights into Field Trends via IEI and RUI
Figure 7: Temporal Distribution of $IEI$ and $RUI$ in Surveys for Recommender Systems and Language Modeling: Higher $IEI$ and $RUI$ values in language modeling over the past two years indicate a trend of intensified development. Although surveys in popular fields tend to attract more academic attention, they also require authors to make greater efforts to keep the content up-to-date.
In Fig. 7 (a), we observe a rapid growth trend in the development of surveys within the field of “language modeling” since 2023, which is reflected by the increasing number of surveys with positive $IEI$ values. Fig. 7 (b) further supports this finding. Since 2023, the median $RUI$ for language modeling surveys has noticeably exceeded that of surveys in the recommendation systems field. As $RUI$ is positively correlated with the volume of publications within a field, a higher $RUI$ suggests a larger publication volume. While this trend underscores the swift progress in the language modeling domain, it also highlights the growing burden on authors to produce and maintain high-quality surveys in such a rapidly evolving and highly competitive area.
By analyzing the statistical characteristics of numerous review articles’ IEI and RUI, we can uncover developmental patterns across various fields. These patterns provide valuable insights for gauging the momentum of progress within each field and advancing exploration in emerging areas.
5 Indicators-Driven Approach for Screening Reviews
@@
As the number of published reviews continues to grow, it has become common for multiple reviews to cover similar areas, attempt to answer analogous research questions, and often arrive at comparable conclusions. This overlap creates information redundancy and, to some extent, reduces research efficiency. To enhance research efficiency, scholars often rely on academic search engines such as Google Scholar, Semantic Scholar, and Web of Science. They tend to prioritize high-ranking results, typically sorted by citation counts, to identify a select number of highly relevant review articles for detailed examination and analysis. However, these results are sometimes perceived as potentially biased or manipulated [47, 76]. Additionally, an over-reliance on citation counts as a metric can intensify the Matthew effect, which may ultimately hinder sustainable academic development within the field. A feasible alternative is to draw on a set of evaluation indicators rather than relying exclusively on citation counts to inform the selection process.

### v6

@@ 2.3.3 Meta-data Statistics of RiPAMI
Figure 3: Statistics of the RiPAMI database: The samples in the RiPAMI database are characterized by a diversity of publication dates, scholar impact, reference numbers, etc., offering a comprehensive reflection of the state of reviews in the PAMI field.
@@
Number of the Authors A review paper typically aims to provide a comprehensive overview of a specific topic. Involving multiple authors with diverse expertise could enhance the depth and breadth of the review. Fig. 3 (b) indicates that the majority of reviews are written by fewer than 10 authors.
Number of the References The number of references in a survey paper may influence its credibility and reliability. As shown in 3 (c), the distribution of literature review references in the RiPAMI database follows a log-normal pattern.
Number of the Citations We count the citations of papers in RiPAMI and plot them in Fig. 3 (d). A power law distribution of received citations could be found. This phenomenon where a small subset of papers receives the majority of citations is sometimes referred to as the ``Matthew Effect" or the ``Pareto principle", as reported in [85, 18].
Length of Literature Reviews The words count and pages number of a review article can partially reflect its depth and breadth. In general, more detailed and comprehensive reviews tend to contain a higher word count and longer pages, thus requiring more time and effort to conduct and comprehend. As shown in Fig. 3 (e) and (f), review articles in the PAMI field typically range from 5,000 to 20,000 words and span 10 to 40 pages. At an average reading speed of 240 words per minute [17], reading a full review would take over 20 minutes.
@@
Further statistical details, such as maximum (Max), minimum (Min), mean, median, and mode, are available in Tab. 2.
Table 2: Statistical Meta-data Summary of RiPAMI
@@
2.4 Bibliometric Indicators for Reviews
Literature reviews reflect the state of their respective fields, highlighting current research priorities and offering a glimpse into emerging trends. Despite their significance, systematic methods for evaluating literature reviews remain underdeveloped. By introducing quantitative tools, bibliometric methods offer a promising solution, enabling researchers to objectively analyze literature reviews and uncover key patterns and trends. Bibliometrics is a research field focused on the use of quantitative methods and statistical analysis to evaluate various aspects of scholarly publications [15]. These methods and analysis, including widely recognized metrics like citation counts and impact factors, offer researchers valuable insights in their daily scientific work, aiding them in making prompt and informed decisions. However, most existing bibliometric methods suffer from several limitations including unfair comparison, misuse, and manipulation as introduced in the ``San Francisco Declaration on Research Assessment'' (DORA) [20] and the Leiden Manifesto [42]. Notably, many of these metrics were not specifically designed for evaluating literature reviews, further limiting their applicability in this context.
2.4.1 Topic Normalized Citation Success Index (TNCSI)
To fairly compare the external influence of literature reviews across different fields within an acceptable cost, we propose guiding Large Language Models (LLMs) to generate the topic key phrase and then calculating the success index of the current review within the same topic (see Appendix for details). The $TNCSI$ estimates the impact of research publications within an LLM-generated topic by normalizing citation counts to a scale between 0 and 1. The formula for calculating the $TNCSI$ is:
$TNCSI=\int_{0}^{\text{citeNum}}\lambda e^{-\lambda x}\,dx,x\geq 0,$ (1)
where $\lambda$ is the scale parameter obtained through maximum likelihood estimation. Detailed steps for estimating $\lambda$ and constructing the probability density function are provided in A.1.
The $TNCSI$ demonstrates favorable mathematical properties and interpretability. First, the $TNCSI$ algorithm employs maximum likelihood estimation to convert the probability mass function into a probability density function. This process ensures that, in theory, the $TNCSI$ differentiates between papers with distinct citation counts, avoiding the assignment of identical values to them. Second, the $TNCSI$ possesses physical significance, representing the probability that a specific paper's citation count surpasses that of any other paper on the same topic. For example, a paper with a $TNCSI$ of 0.5 means it has more citations than half of the papers within the same topic. Furthermore, the calculation of $TNCSI$ is computationally efficient, as it avoids requiring a complete or precise ranking of all related papers. Instead, it provides a reliable estimation of the probability that the given paper surpasses others in terms of citation count within the same topic.
2.4.2 Impact Evolution Index (IEI)
The $IEI$ is defined as the average slope across $l$ distinct points on a n=l1n=l-1 degree Bézier curve representing the citation trend over time. The formula for calculating $IEI_{L_{l}}$ is:
$IEI_{L_{l}}=\sum_{a=0}^{l-1}\frac{(y_{i}/x_{i})}{l},$ (2)
where $l$ represents the number of months observed. $x_{i}$ and $y_{i}$ are the components of the tangent vector at the $a$-th point on the Bézier curve, indicating the magnitude along the $x$- and $y$-axes, respectively. More details can be found in A.2.
2.4.3 Reference Quality Measurement (RQM)
$RQM=1-e^{-\beta\cdot e^{-(1-ARQ)\cdot S_{mp}}}.$ (3)
In this equation, $\beta$ represents the shift parameter, $ARQ$ denotes the average reference quality, while $S_{mp}$ refers to the median semester count of the reference age, indicating the time from the publication date of the references to the issuance of the review (which will be detailed in  A.3.
Beyond simply assessing the average impact of references, the $RQM$ integrates references' timeliness and its influence on review quality. This metric addresses the tendency among authors to rely heavily on classic works, which may no longer fully align with current research trends. Additionally, the adjustment coefficient in $RQM$ is calibrated through statistical and optimization algorithms, tailored to the unique characteristics of each research discipline. This data-driven approach minimizes reliance on heuristic parameter settings and increases RQM's adaptability across various academic disciplines.
2.4.4 Review Update Index (RUI)
$RUI=p\cdot CDR+q\cdot RAD.$ (4)
Here, $p$ and $q$ are weighting coefficients, set to 10 and 5, respectively. The 2:1 ratio ensures that $CDR$ plays a dominant role, while the specific magnitudes primarily enhance numerical readability rather than altering the substantive balance between the terms. Detailed definitions and calculations of the $CDR$ and $RAD$ are provided in the A
2.5 Information Extraction for Reviews
Given that this study aims to identify the common features of reviews within the field, solely relying on manual inspection would be both impractical and resource-intensive. To further investigate the features of reviews in the PAMI field, we employed LLM-based information extraction techniques to analyze the review content and parse the response to structured results. The proposed approach converts unstructured text into structured data, facilitating further analysis and interpretation. Details regarding the information extraction method can be found in the  B.
3 Review of Reviews in PAMI
This section provides an integrated overview of the PAMI field by combining narrative synthesis and statistical analysis, aiming to identify common features of existing reviews and to uncover quantitative patterns across various sub-fields, thereby laying the foundation for subsequent in-depth discussion.
3.1 Narrative Reviews in Various Sub-fields in PAMI
In this subsection, we offer a subjective analysis of reviews from diverse fields. Through this narrative approach, we aim to examine the composing characteristics shared by reviews across various fields. It should be noted that only a subset of fields is discussed here for illustrative purposes, while the complete list of covered sub-fields is provided in D
3.1.1 Computer Vision
3.1.2 Natural Language Processing
Sentiment Analysis focuses on determining the sentiment or emotion expressed in text, such as positive, negative, or neutral. Yadav et al. introduce the process of gathering and analyzing people's opinions and sentiments from various sources such as social media platforms and blogs in their paper [140]. The paper evaluates and compares different approaches used in sentiment analysis, with a focus on supervised machine learning methods like Naive Bayes and SVM algorithms. The common application areas of sentiment analysis and the challenges involved in accurately interpreting sentiments are also reported. The paper by Yue [147] categorizes and compares a large number of techniques and methods from three different perspectives: task-oriented, granularity-oriented, and methodology-oriented. It also explores different types of data and advanced tools for research, highlighting their strengths and limitations.
3.1.3 Others
3.2 Popular Structure of Reviews in PAMI
The review's structure is alternatively referred to as the framework of the review. It is the outline that authors consider when starting to conduct the survey. Generally, a well-designed structure is considered essential for enhancing a paper’s readability and helping readers grasp its core concepts and knowledge. Such a framework typically includes foundational components such as the abstract, introduction, methodology, discussion, and conclusion, each serving a distinct communicative function. To investigate how these elements are actually organized, we conducted a large-scale analysis of section titles across our collected corpus of survey papers. We first identified a representative set of section keywords (e.g., Introduction, Methods, Experiments, Results, Discussion, Future Work, Conclusion) and extracted their occurrences. For each keyword, we normalized its position within the document (ranging from 0 at the beginning to 1 at the end) and then aggregated these values across all papers. The resulting distributions, illustrated with violin plots in Fig. 4, reveal clear structural anchors: ``Introduction'' consistently appears at the very beginning, while ``Conclusion'' is concentrated at the end. Expository sections such as ``Background'' and ``Preliminaries'' occur early; ``Related Work'' and ``Taxonomy'' appear in the early-to-middle part; and ``Methods'' and ``Dataset'' are most often located around the middle. Empirical sections (``Experiments'', ``Results'', ``Evaluation'') tend to cluster in the latter half, whereas reflective components (``Application'', ``Discussion'', ``Challenges'', ``Future'') are predominantly placed near the end. More detailed information on the algorithm can be found in B.4.
Figure 4: Distributions of normalized section positions across survey papers: Core sections such as Introduction and Conclusion exhibit highly consistent placement at the beginning and end, while methodological, empirical, and reflective components show broader variation across the document body.
@@ 3.2 Popular Structure of Reviews in PAMI
To gain a deeper understanding of structural practices, we applied information extraction techniques to the RiPAMI database. Using an ensemble query approach based on LLMs, we automatically identified six representative content features: (1) Taxonomy — whether the review explicitly outlines a taxonomy of methods; (2) PRISMA — whether it specifies inclusion/exclusion criteria akin to PRISMA guidelines [94]; (3) Preliminary — whether it provides background knowledge in a distinct section; (4) Benchmark — whether it presents quantitative comparisons across methods; (5) Application — whether it discusses real-world applications; and (6) Discussion — whether it addresses challenges, limitations, or future directions. Each feature was recorded in binary format (0/1) for statistical analysis (see B for extraction details).
The left panel of Figure 5 presents the overall prevalence of the six identified features. A majority of reviews include a dedicated discussion section on challenges and future directions, reflecting the academic expectation that surveys should not only summarize existing work but also identify open problems and potential avenues for research. In contrast, fewer than 20% of reviews in the PAMI field adopt PRISMA-like methodological criteria, regardless of publication venue, revealing substantial room for more rigorous systematic practices. The right panel of Figure 5 highlights domain-specific differences: reviews in computer vision (CV) show a pronounced emphasis on benchmarking, whereas those in smaller or cross-disciplinary areas (MISC) more frequently incorporate preliminaries.
Figure 6 illustrates temporal trends. All six features have increased steadily over the past decade, with the inclusion of Discussion'' and Preliminary'' sections showing the sharpest rise. Particularly noteworthy is that the adoption of PRISMA-style reporting has grown by more than 10% between 2015 and 2024, pointing to a gradual shift toward more structured and methodologically robust reviews within the PAMI community.
@@
Figure 5: Statistical Analysis of Review Features in the RiPAMI Database using LLM-based Information Extraction. “Published” refers to peer-reviewed review papers, whereas “Non-published” denotes preprints without peer review. CV, NLP, and MISC correspond to the fields of computer vision, natural language processing, and miscellaneous domains, respectively.
@@
Building on the preceding analysis, we now discuss the four major components commonly observed in survey papers: the abstract, the beginning part, the middle part, and the ending part.
3.2.1 Abstract
The abstract is widely regarded as the most important part of a paper, as it provides a concise summary of the core content, enabling readers to quickly grasp the main points and conclusions. While the level of detail may vary, a structured abstract improves clarity and reproducibility by dividing the content into clearly labeled sections, each fulfilling a specific communicative function. Importantly, international reporting standards such as the PRISMA guidelines [94] explicitly recommend the use of structured abstracts in systematic reviews and meta-analyses, as they enhance transparency and facilitate information retrieval [56, 12]. A typical structured abstract consists of sections such as Background, Objective, Methods, Results, Conclusion/Discussion, and sometimes Limitations. For instance, Sariyanidi et al. [104] structured their abstract by introducing the background on facial affect, highlighting persistent challenges in the field, and then summarizing their methods, results, and discussion.
We further instruct the LLM to assess whether an abstract adheres to the structured format outlined by PRISMA guidelines (detailed in  B). As shown in Fig. 7 (a), the proportion of structured abstracts in the PAMI field remains relatively low, suggesting that this field has yet to fully embrace the structured abstract format. However, panel (b) illustrates a clear upward trend in the number of abstracts meeting the structured abstract requirements, with a steady increase observed year on year. This highlights a growing adoption of standardized abstract formats in recent research.
Figure 7: Proportion and Trend of Structured Abstracts: The bar chart on the left shows the proportion of structured abstracts across different fields (CV, NLP, MISC), while the line chart on the right illustrates the trend in the number of structured abstracts over the years from 2010 to 2024.
3.2.2 The Beginning Part
Similar to the research article, the introduction section of the literature review usually includes contextualizing the research topic, identifying the knowledge gap, defining the scope and objectives, and laying the foundation knowledge for wider audiences. ``Introduction'' usually lies at the very beginning of the paper. Most reviews first introduce the definition of the topic to acquaint readers with a basic understanding of the field. For instance, the term ``Named Entity Recognition'' may be unfamiliar to many scholars outside the field of natural language processing. A clever introduction might provide the origins of the terminology and inform the readers what is named entity recognition, as has been done in Ref [90]. A brief definition of the field in the introduction is also welcomed for some relatively popular fields, e.g., the literature review [103] examines the concept of image categorization in the first sentence.
Preliminaries and problem formulations are also popular with readers, as they both provide profound background knowledge. Given that the Gumbel-max involves considerable mathematical concepts and calculations, a ``Preliminaries'' section in the review of the Gumbel-max trick [45] serves the purpose of providing background information and basic understanding related to the Gumbel-max trick.
For systematic reviews, a clear statement regarding the review methodology (e.g., literature search strategy and criteria for inclusion and exclusion) is essential [94, 55]. Such a statement helps to minimize potential bias and enables other researchers to replicate the study's findings. Memon et. at. [84] dedicate a section to detailing their review process, including subsections on the review protocol, inclusion and exclusion criteria, search strategy, study selection process, quality assessment criteria, and data extraction and synthesis, ensuring transparency and rigor in their methodology.
3.2.3 The Middle Part
The middle part of a survey paper, also known as the review part, presents a detailed examination of relevant research studies, methodologies, findings, and theories related to the chosen theme. Beyond mere description, it synthesizes information from the literature to provide a cohesive and integrated understanding of the topic. This synthesis not only aligns disparate works but also evaluates their contributions and interrelations. It often includes comparative analysis, highlighting similarities and differences in approaches, and may identify limitations in the existing research that warrant further investigation.
Across existing surveys in the PAMI field, different organizational strategies can be observed. Some reviews arrange the discussion by grouping methods according to their technical characteristics, thereby creating typologies that help readers compare families of approaches. For instance, Minaee et al. conducted a comprehensive survey on deep learning-based image segmentation models [87], grouping them into ten categories such as fully convolutional models, encoder–decoder-based models, and attention-based models. This mode of presentation highlights the breadth of methodological options, but can be less accessible for readers who are not yet familiar with the field.
Other surveys are organized around specific research problems or tasks, with each section devoted to a challenge and the solutions proposed for it. A review on egocentric vision hand analysis by Bandini et al. [8] illustrates this approach: the paper is divided into sections such as hand segmentation, hand detection, and hand identification, with subsections that further explore issues like robustness to illumination changes or lack of pixel-level annotations. This problem-driven framing can be especially useful in application-oriented areas, though it may offer less systematic elaboration of theoretical underpinnings.
In many cases, authors blend these strategies. A review may be structured by tasks or challenges, while within each task methods are further grouped and compared. Guo et al. [38], for example, divided their survey into task-based sections but introduced similar approaches group by group within each task. Such a blended organization can provide both problem-driven context and methodological coverage, giving readers multiple perspectives on the same body of work.
Overall, the middle part of a survey is where synthesis and comparison take place, and where authors draw together diverse contributions into a structured narrative. Whether arranged around methods, challenges, or a combination of both, these organizational choices illustrate how surveys strive to balance comprehensiveness, accessibility, and clarity in guiding readers through complex literature.
3.2.4 The Ending Part
The subsequent ending part serves as a succinct conclusion to the reviewed studies. Within this section, the authors prefer to sum up the key findings to answer the question in the beginning part. It usually highlights both the advantages and disadvantages of the reviewed literature to conclude research gaps and suggest potential avenues for future research.
One of the main objectives of the concluding part is to provide a concise summary of the key insights derived from the reviewed studies and to emphasize their significance and relevance to the research topic. In some literature reviews, however, the analysis and synthesis are so extensive that the main body spans a considerable number of pages. For instance, Liu et al. [65] present a survey whose main body covers 51 double-column pages, posing a challenge for readers who may not have the time to read it thoroughly. Fortunately, the paper also provides a concise summary, which enables readers to quickly grasp the essential content and navigate to sections of particular interest.
When discussing further in the ending part, the authors may attempt to identify gaps and point out future directions through the analysis of existing literature. Some of the papers present the gaps and future directions separately. Hassain et al. [46] first discussed major challenges of multi-view video summarization such as lack of synchronization, instability of camera, and crowded scenes individually. Where there are challenges, there are future research directions. In addition to challenges, the authors also provide recommendations and future directions from various perspectives including models, benchmark datasets, and agents-based MVS, etc. Conversely, some papers choose to combine discussions of current issues with emerging research trends, as seen in the 'Future Directions' section of Ref[143].
4 Indicator-Guided Navigation for Literature Reviews
@@
As the number of published reviews continues to grow, it has become common for multiple surveys to cover similar areas, attempt to answer analogous research questions, and often arrive at comparable conclusions. This overlap creates information redundancy and, to some extent, reduces research efficiency. Human scholars frequently rely on academic search engines such as Google Scholar, Semantic Scholar, and Web of Science, prioritizing high-ranking results typically sorted by citation counts to identify a select number of reviews for closer examination. However, these rankings are sometimes perceived as biased or even manipulated [48, 78], and an over-reliance on citation counts can intensify the Matthew effect, which may ultimately hinder sustainable academic development within the field.
Building on the indicators introduced in Sec. 2, this section demonstrates how they can be applied to enable indicator-guided navigation. Such navigation not only helps researchers more effectively evaluate and select reviews, but also provides a principled mechanism for AI-for-Research systems that use human-authored surveys as starting points for automated outline generation and knowledge synthesis. In this way, the indicators serve as a bridge between methodological rigor and practical strategies for navigating the expanding corpus of reviews.